### v1

# Reinforcement Learning Pair Trading: A Dynamic Scaling Approach

Abstract
Cryptocurrency is a cryptography-based digital asset with extremely volatile prices. Around \$70 billion worth of crypto-currency is traded daily on exchanges. Trading crypto-currency is difficult due to the inherent volatility of the crypto-market. In this work we want to test the hypothesis: “can techniques from artificial intelligence help with algorithmically trading crypto-currencies”. In order to address this question; we combine Reinforcement Learning (RL) with pair trading. Pair trading is an statistical arbitrage trading technique, which exploits the price difference between statistically correlated assets. We train reinforcement learners to determine when and how to trade pairs of crypto-currencies. We develop new reward shaping and observation/action spaces for reinforcement learning. We performed experiments with the developed reinforcement learner on pairs of BTC-GBP and BTC-EUR data separated by 1-minute intervals ($n=263,520$). The traditional non-RL pair trading technique achieved annualised profit of 8.33%, while the proposed RL based pair trading technique achieved annualised profits from 9.94% — 31.53%, depending upon the RL learner. Our results show that RL can significantly, outperform manual and traditional pair trading techniques when applied to volatile markets such as crypto-currencies.
keywords:
Pair Trading , Reinforcement Learning , Algorithmic Trading , Deep Learning , Cryptocurrency
journal: Expert Systems with Applications
1 Introduction
Arbitrage is a subdomain of financial trading that profits from price discrepancies in different markets [Dybvig & Ross, 1989]. Pair trading is one of the well-known arbitrage trading methods in financial markets. Arbitrageurs identify two highly correlated assets to form a pair. When a price discrepancy happens, they buy the underpriced asset and sell the overpriced correlated asset to profit from the mean reversion of the prices. Arbitrage requires practitioners to constantly analyse the market conditions at the fastest speed possible, as arbitrageurs must compete for transitory opportunities [Brogaard et al., 2014]. The faster the market analysis is, the more the chances of profiting from arbitrage. Therefore, we want to explore the process of utilising Artificial Intelligence (AI)to accelerate the process of pair trading.
Reinforcement Learning (RL)is a captivating domain of AI. The idea of RL is to let the agent(s) learn to interact with an environment. The agent should learn from the environment’s responses to optimise its behaviour [Sutton & Barto, 2018]. If we view the financial market from the perspective of RL environment, actions in the financial market are investment decisions. To gain profits, arbitrageurs are incentivized to train agents that produce lucrative investment decisions, and RL facilitates agents’ learning process from the profit/loss of the market.
The combination of RL and various financial trading techniques is still going through rapid evolution. There have been some work in RL infrastructural construction [Liu et al., 2021, 2022b, 2022a], and some experiments in profitable RL agent training [Meng & Khushi, 2019, Zhang et al., 2020, Pricope, 2021]. Trading actions in traditional pair trading follow static rules. In reality, the complexity of financial markets should allow more flexibility in the decision-making process. An experienced trader might analyse the market conditions to make informed decisions. However, it is not feasible to output efficient decisions at short, intermittent intervals 24/7. RL algorithms enable a fast-track decision-making process for analysing trading signals and generating trading actions.
In designing a high-frequency trading system based on RL, several problems must be explored to ensure a fast decision-making process. The first problem is pair formation: identifying compatible instruments with historical correlations to form profitable pairs. The second problem concerns timing: instead of blindly following preset rules, the system needs flexibility in choosing investment timing for greater profit. The last challenge involves investment quantity: as investment opportunities vary in quality, experienced traders can select better opportunities with stronger profit potential. It is worth investigating whether RL agent is capable of achieving similar profitability by scrutinizing each investment opportunity.
This paper investigates some questions centred around RL in pair trading. To overcome the fast decision-making requirement in a high-frequency trading environment, we constructed an environment that suits the RL agent to conduct pair trading and fine-tune reward shaping to encourage the agent to make profitable decisions. The contributions of this paper are: (i) proposal of a novel pair trading method that is adaptive to high volatility markets. (ii) utilisation of grid search technique to fine-tune profitable hyperparameters in pair trading. (iii) introduction of the RL component in pair trading for market analysis and decision-making. (iv) development of a novel RL model for making decisions about the quantity of investment.
The structure of the paper is arranged as follows: the background and related work are introduced in Sections 2 and 3. The methodology is presented in Section 4. Experiments and results are included in Section 5. Following by discussion about the results and conclusion in Section 6.
2 Background
First, we define the basic terms of financial trading. A long position is created when an investor uses cash to buy an asset, and a short position is created when an investor sells a borrowed asset. The portfolio is the investor’s total holding, including long/short position and cash. Transaction cost is a percentage fee payable to the broker for any long/short actions. Finally, risk is defined as the volatility of the portfolio.
2.1 Traditional Pair Trading
Classical pair trading consists of two distinct components known as legs. a leg represents one side of a trade in a multi-contract trading strategy. Under the definition of pair trading, “longing the first asset and shorting the second asset” is called long leg, and “shorting the first asset and longing the second asset” is called short leg. The two assets are always bought and sold in opposite directions in pair trading. Therefore, the overall pair trading strategy is considered to be market neutral, because the profits from the long position and the short are offset by the direction of the overall market. Gatev et al.’s 2006 work is the most cited traditional pair trading method. It follows the OODA (Observe, Orient, Decide and Act) Loop [Fadok et al., 1995]. Before entering the market, the first step is to choose the proper assets in a pair. Sum of Squared Deviation (SSD)is the measurement calculated from prices for assets $i$ and $j$. Through exhaustive searching in a formation period $T$, the assets with the smallest SSD are bound as a pair (Equation 1).
$SSD_{p_{i},p_{j}}=\sum_{t=1}^{T}(p_{i}-p_{j})^{2}$ (1)
  • 1.
    Observe is the process of market analysis. The price of assets in pairs is collected and processed. The price difference $(p_{i}-p_{j})$ is called Spread $S$. The arbitrageurs observe the current positions and spread of the current market.
  • 2.
    Orient is exploring what could be done. Three possible actions for pair trading are long leg, short leg and close position as defined above.
  • 3.
    Decide what action to take. Position opening triggers when the price difference deviates too much. This is indicated by the spread movement beyond an open threshold. Position closing happens when the spread reverts back to some closing threshold. Gatev et al. (2006) adopted two times the standard deviation of the spread as the opening threshold and the price crossing as the closing threshold. In practice, the threshold varies according to the characteristics of the financial instrument.
  • 4.
    Act once the decision is made. The long leg orders to buy asset $i$ and sell asset $j$; The short leg orders to sell asset $i$ and buy asset $j$. Closing a position means clearing all the active positions to hold cash only.
Figure 1: Price distance view of pair trading $p_{i}$ and $p_{j}$
A graphical visualisation of pair trading is presented in Figure 1. Figure 1 (a) is the market interactions according to the spread and thresholds. A position is opened whenever the spread deviates beyond the open threshold. The position closure happens when the spread reverts below the close threshold. Figure 1 (b) presents the corresponding actions with the degree of spread deviation. The spread deviations are classified into zones based on the Spread ($S$), Open-Threshold ($OT$) and Close-Threshold ($CT$) (Equation 2):
$\left\{\begin{array}[]{rl}\text{Short Zone:}&+OT<S<+\infty,\\ &\text{spread deviates beyond open threshold}\\ \text{Neutral Short Zone:}&+CT<S<+OT,\\ &\text{spread deviates between open and close threshold}\\ \text{Close Zone:}&-CT<S<+CT,\\ &\text{spread reverts between close thresholds}\\ \text{Neutral Long Zone:}&-OT<S<-CT,\\ &\text{spread deviates between open and close threshold}\\ \text{Long Zone:}&-\infty<S<-OT,\\ &\text{spread deviates below open threshold}\\ \end{array}\right.$ (2)
2.2 Reinforcement Learning
Reinforcement Learning (RL)is used for training an agent for maximising rewards while interacting with an environment [Sutton & Barto, 2018]. The environment for RL is required to be a Markov Decision Process (MDP[Bellman, 1957], which means it is modelled as a decision-making process with the following elements 〈State ($S$), Action ($A$), Transition ($P_{A}$), Reward ($R_{A}$)⟩. The goal is to train the agent to develop a policy ($\pi$) that fulfil and objective, e.g., maximising profits in a trade. At every trading interval $t$, according to the state $S$, that the agent observes, an action $A$ is chosen based on the policy $\pi$. The environment rewards/punishes the state transition of $S_{t}\rightarrow S_{t+1}$ with environment reward $r$. If we assume $\gamma$ to be the discount factor for the time-value discount of future reward, RL trains a policy $\pi$ that maximises the total discounted reward $G_{t}$ as shown in Equation 3:
$G_{t}=\sum_{i=0}^{\infty}\gamma^{i}r_{t+i}$ (3)
RL algorithms can broadly be classified based on three criteria: value/policy-based, on/off-policy and actor/critic-network [AlMahamid & Grolinger, 2021]. Value-based methods estimate state-action value functions for decision-making. Policy-based methods directly learn action selection policies. The on-policy method requires data generated by the current policy, and off-policy is capable of leveraging past experiences from potentially different policies. Moreover, actor-critic architectures, where an actor-network proposes actions and a critic network evaluates, have shown a better performance in facilitating policy improvement through this feedback loop. Most recent researches favour actor-critic architecture instead of actor-only or critic-only methods for better performance [Meng & Khushi, 2019, Zhang et al., 2020]. Therefore, only actor-critic algorithms are adopted in this study.
Based on the RL classification criteria, some representative algorithms have been selected for this study including Deep Q-Learning (DQN)[Mnih et al., 2013], Soft Actor Critic (SAC)[Haarnoja et al., 2019], Advantage Actor-Critic (A2C)[Sutton & Barto, 2018], Proximal Policy Optimization (PPO)[Schulman et al., 2017]. Diversified RL algorithms are experimented with to choose the most effective one in pair trading.
3 Related Work
3.1 Reinforcement Learning in Algorithmic Trading
Reinforcement Learning in AlphaGo captured the world’s attention since 2016 by participating in a series of machine versus human competitions on board game GO [David Silver, Demis Hassabis, 2016]. Surprisingly, the research regarding RL in the financial market started long before that. Recurrent reinforcement learning studies were the mainstream works [Gold, 2003, Bertoluzzo & Corazza, 2007, Maringer & Ramtohul, 2012, Zhang & Maringer, 2016] in the early stage of financial trading. After the upsurge of AlphaGo, some significant advancements were brought to RL trading as well, Huang re-described Markov Decision Process (MDP)financial market as a game process to incorporate RL as Financial Trading as a Game [Huang, 2018]. Newer RL models such as Deep Q-Learning (DQN), Policy Gradients (PG)and Advantage Actor-Critic (A2C)have also been used recently by researchers [Meng & Khushi, 2019, Zhang et al., 2020, Pricope, 2021] for financial trading. A noteworthy research work is that of FinRL group in the infrastructures and ensemble learning mechanism [Liu et al., 2021, 2022b, 2022a].
3.2 Reinforcement Learning in Pair Trading
Reinforcement Learning, in combination with pair trading, is not an untapped domain. RL has ameliorated multifaceted aspects from the traditional method of pair trading brought up by Gatev et al. 2006. RL technique Ordering Points To Identify The Clustering Structure (OPTICS)contributed at the pair selection stage by leveraging a clustering algorithm to produce better pair choices [Sarmento & Horta, 2020]. Vergara & Kristjanpoller 2024 brought deep reinforcement learning into the Cryptocurrency world in an ensemble setting. reCurrent Reinforcement lEarning methoD for paIrs Trading (CREDIT)algorithm that takes consideration of both profitability and risks was engineered by Han et al. 2023. Reward Shaping is also an interesting area where some work has been done in RL trading [Lucarelli & Borrotti, 2019, Wang et al., 2021]. Kim & Kim’s work 2019 is the most recent RL pair trading method. Their focus is on utilising RL to find most trading opportunities. Instead of fixed thresholds, the RL agent in Kim & Kim’s work produces the thresholds for the coming trading period. Open, close and stop-loss thresholds determine the profits of pair trading.
Our work introduces a novel method to combine RL with pair trading. The work of Gatev et al. is not efficient enough for a high-frequency market. The state-of-the-art method of Kim & Kim has some deficiencies: (i) it requires the market’s volatility to be relatively stable. The RL agent may produce unsuitable thresholds if the market experiences increased volatility. (ii) It lacks flexibility in the investment amount. Opportunities with different qualities are programmed to be invested with the same amount of capital. Once the RL agent determines a threshold, the trading algorithm executes a trade at pre-determined thresholds. We leverage RL to make investment timing and quantity decisions. The adjustable investment amount is a novel feature of our RL pair trading. RL agent measures how well the investment opportunities are based on observations and invests a larger amount in more promising market conditions. Having another dimension on the investment side should further enhance profitability and reduce risks.
4 Methodology
In this section, we introduce the architecture of the methodology (Figure 2). The architecture includes five steps: (1) Pair Formation for selecting assets to form a tradeable pair (Section 4.1); (2)Spread Calculation utilising the moving-window technique to extract the spread in a limited retrospective time frame (Section 4.2); (3) Parameter Selection from historical dataset to decide the most suitable hyperparameters for pair trading (Section 4.2); (4) RL Trading by allowing RL to decide the trading timing and quantity in pair trading (Section 4.4); (5) Investment Action for taking the actions produced from RL trading into market execution.
Figure 2: Architecture of Trading Strategies.
4.1 Pair Formation
Pairs are selected based on two criteria — correlation and cointegration. The widely adopted Pearson’s correlation [Perlin, 2007, Do & Faff, 2010] is given by
$\rho_{X,Y}=\frac{\mathrm{cov}(X,Y)}{\sigma_{X}\sigma_{Y}}$ (4)
where $\rho_{X,Y}$ is the correlation coefficient between assets $X$ and $Y$, $\mathrm{cov}(X,Y)$ is the covariance of $X$ and $Y$, and $\sigma_{X}$ and $\sigma_{Y}$ are the standard deviations of $X$ and $Y$, respectively. The Engle-Granger cointegration test [Burgess, 2003, Dunis & Ho, 2005] involves two steps. First, the linear regression is performed:
$Y_{t}=\alpha+\beta X_{t}+\epsilon_{t}$ (5)
where $Y_{t}$ and $X_{t}$ are the asset price series, $\alpha$ and $\beta$ are the regression coefficients, and $\epsilon_{t}$ is the residual term. The second step tests the residuals $\epsilon_{t}$ for stationarity using an Augmented Dickey-Fuller (ADF)[Dickey & Fuller, 1979] test. The ADF test regression is given in Equation 6.
$\Delta\epsilon_{t}=\gamma\epsilon_{t-1}+\sum_{i=1}^{p}\delta_{i}\Delta\epsilon_{t-i}+\nu_{t}$ (6)
where $\Delta\epsilon_{t}$ is the first difference of the residuals, $\gamma$ is the coefficient to be tested for stationarity, $p$ is the number of lagged difference terms included, and $\nu_{t}$ is the error term. If $\gamma$ is significantly different from zero, the residuals are stationary, indicating cointegration.
A moving window is slid over historical pricing data (Figure 3). Averaged correlation and cointegration batches are used in the selection phase to ensure that the assets selected have a strong long-term statistical relationship.
Figure 3: Window-size cut for correlation and cointegration test
4.2 Spread Calculation
The second step is a moving window mechanism to capture the spread movement (Figure 2 Step-2). Spread $\epsilon_{t}$ is calculated at every selected trading interval (e.g., every five minutes) is the error term $s$ from a regression between the two prices $p_{i}$ and $p_{j}$, which follows normal distribution with mean of 0 and standard deviation of $\sigma$ (Equation 7):
$p_{i}=\beta_{0}+\beta_{1}\times p_{j}+s_{i}\sim N(0,\sigma^{2})$ (7)
We normalise the spread with $z$-score (Equation 8) to scale the spread into constant mean and standard deviation. The mean of the spread in the sliding window is represented as $\bar{s}$:
$Z=\frac{{s-\bar{s}}}{{\sigma_{s}}}$ (8)
4.3 Parameter Selection
Gatev et al. [2006] adopted 2 times standard deviation as the open threshold and the deviation crossing point as the close threshold (Figure 1). However, the thresholds ought to vary with market conditions. Hence, a window-sliding mechanism is incorporated to reflect the heterogeneity of the pricing variance [Mandelbrot, 1967].
Three parameters to be explored are 〈Window Size, Open Threshold, Close Threshold⟩. Window size is the number of historical samples in the moving window. Thresholds are highly linked to market conditions. Excessively wide thresholds suit more volatile markets, and conservatively narrow thresholds show result in smaller but steadier wins. The combination of parameters of the highest profitability 〈Window Size, Open Threshold, Close Threshold⟩ are selected from a search pool through grid search in practice.
4.4 Reinforcement Learning Pair Trading
After we run the grid search of window-sliding pair trading, the next problem concerns “when” and “how much” to trade. The pair trading results from following pre-set rules (à la Gatev et.al. 2006) are obtainable using window-sliding pair trading. However, we want to know if RL produces better investment decisions than blindly following the rules. Therefore, the most profitable parameter combination is passed onto further RL-based pair trading so that we can compare the results between RL-based pair trading and non-RL pair trading.
4.4.1 Observation Space
Observation space stands for the information an RL agent observes. The agent observes the market information to make decisions. The observations adopted for our RL environment is the following tuple: 〈Position, Spread, Zone⟩.
  • 1.
    Position $\in[-1,1]$: Position stands for the current portfolio value. Position is a percentage measuring the direction of investment (c.f. Figure 4). Assuming that we do not use leverage, holding a long leg with 70% portfolio value gives Position = 0.7. Holding a short leg with 30% portfolio value gives Position = -0.3. Position 0 means we only hold cash.
    Figure 4: The value of position observation based on investment
  • 2.
    Spread $\in\mathbb{R}$: It represents how much the current spread has deviated from the mean (Section 4.2).
  • 3.
    Zone $\in\{\text{Zones}\}$: Zone is an important indicator that comes from the comparison between the $z$-score with the thresholds for signals (Figure 1(b)). Traditional pair trading [Gatev et al., 2006, Yang & Malik, 2024] takes the zone as the direct trading signal (Figure 1(c)). However, in RL-based pair trading, the zone is an observation for the RL agent to make better decisions.
4.4.2 Action Space
Since the pair trading technique is a relatively low-risk strategy, most applications invest with a fixed amount or the complete portfolio value [Burgess, 2003, Perlin, 2009, Huck, 2010]. Meanwhile, it is natural for an experienced trader to invest different amounts based on the quality of opportunities. Opportunities with a higher probability of success are worth more investment. Therefore, we investigate granting the RL agent not only the decision of when to invest but also the freedom to choose the investment amount.
Take $A\in[-1,1]$ as the action. Similar to the observation space (c.f. Figure 4), action ranges from -1 to 1, representing the investment as a percentage of portfolio value to the long leg and short leg direction. Investing 50% of the portfolio as a long leg means A = 0.5. Investing 20% of the portfolio in short leg means A= -0.2.
In practice, we have to consider the relationship between the existing position and the next action. We classify the execution of action as 〈Open Position, Adjust Position, Close Position⟩.
  • 1.
    Open Position is the action of opening a new position.
  • 2.
    Close Position is the closure of a position.
  • 3.
    Adjust Position happens when a previous position is open and the RL agent is wants to open another potion. For example, if the current position is 70% long-leg, and the new action is A=0.8, only the extra 10% shall be actioned.
4.4.3 Reward Shaping
The RL reward consists of three components: 〈Action Reward, Portfolio Reward, Transaction Punishment⟩.
  • 1.
    Portfolio Reward is the profit/loss from closing a position. The portfolio value $V_{p}$ updates only on position closing. If $V^{\prime}_{p}$ is the position value at the start of a trading period $p$. Then, upon closing the trade at the end of trading period $p$, the reward it calculated as shown in Equation (9).
    Profit $\text{Profit Reward}=V_{p}-V^{\prime}_{p}$ (9)
  • 2.
    Action Reward means the agent needs to be rewarded for taking a desired action in the corresponding zone. In general, the agent is free to decide on any action. However, we use Action Reward to encourage the agent to choose desired actions with Action Reward. It rewards the agent for making the desired action in certain zones (Table 1) with some freedom in neutral zones. The stronger the action reward, the more it resembles traditional pair trading.
    Zones Rewarding Behaviour
    Short Zone Short leg
    Neutral Short Zone Short leg or Close
    Close Zone Close
    Neutral Long Zone Long leg or Close
    Long Zone Long leg
    Table 1: Rewarding behaviours in zones
  • 3.
    Transaction Punishment is a negative reward for encouraging small adjustments instead of large changes in the position. The punishment is the difference between the action and position. If the current position in observation is $P$ and the action is $A$, the transaction punishment is (Equation 10):
    Transaction $\text{Transaction Punishment}=P-A$ (10)
5 Benchmark Results
Next, we carry out experiments using the proposed methodology. We adopt the same dataset and the same parameters for non-RL pair trading and RL pair trading for comparison purposes.
5.1 Experimental Setup
We experiment with window-sliding pair trading and RL pair trading in the cryptocurrency market. The cryptocurrency market is famous for its volatility, easy access and 24/7 operating time.
5.1.1 Datasets
Figure 5: Prices of BTCEUR and BTCGBP
The application of our trading methodology is on Binance — the largest cryptocurrency market [^1] https://www.binance.com/en. For the best market liquidity, we picked Bitcoin - Fiat currencies under different trading intervals for pair trading. Pair formation criteria are based on Pearson’s correlation and augmented Engle-Granger two-step cointegration test (Section 4.1) for quote currencies that follow a similar trend against the base currency (Figure 5). The formation period is Oct-2023 to Nov-2023, and the test is in Dec-2023, with trading intervals of 1min (121,500 entries), 3min (40,500 entries), and 5min (24,300 entries), respectively. We exhaustively compared correlation and cointegration for the best pair (Table 2)  [^2] While calculating the Cointegration and Correlation, intervals with low volume trades are exempted from the calculation.. Though Binance has quite a few fiat currencies, only the US Dollar (USD), Great Britain Pounds (GBP), Euro (EUR) and Russian Ruble (RUB) display relatively strong liquidity. The pair with the strongest correlation and cointegration is BTCEUR and BTCGBP under 1min trading interval (2).
Pairs 1m 3m 5m
coint corr coint corr coint corr
BTCEUR-BTCGBP 0.5667 0.8758 0.4667 0.8759 0.4667 0.8754
BTCEUR-BTCRUB 0.3333 0.8417 0.3333 0.8417 0.3167 0.8416
BTCEUR-BTCUSD 0.1667 0.9328 0.2000 0.9327 0.2000 0.9329
BTCGBP-BTCRUB 0.3500 0.7606 0.3333 0.7608 0.3333 0.7603
BTCGBP-BTCUSD 0.4833 0.8404 0.4167 0.8403 0.4000 0.8403
BTCRUB-BTCUSD 0.4000 0.8538 0.3333 0.8539 0.3500 0.8543
Table 2: Correlation and cointegration of pair formation
Transaction cost in the experiment is set to 0.02% commission based on Binance’s fee scheme  [^3] https://www.binance.com/en/fee/futureFee. The transaction cost of 0.02% is a flat percentage charge for transactions in both directions. A pair trading leg, including long the first asset and short the second asset, is charged for both long and short actions.
5.1.2 Grid Search and Reinforcement Learning
Grid search is used to find a set of profitable parameters, including open/close threshold and window size during the training period (Oct-2023 to Nov-2023). Every iteration of the window-sliding pair trading experiments with one set of parameters (window size, open/close threshold) until exhaustion. The most profitable parameter set will be used to test traditional pair trading during the test period (Dec-2023) and also for testing the proposed RL-strategy.
The profitability in grid search is measured by Total Compound Return (RTOT), where $V_{p}$ and $V_{p}^{\prime}$ are the value of the portfolio at the beginning of the period and the end of the period and $t$ as the total length of the trading period (Equation 11):
$rtot=(V_{p}^{\prime}/V_{p})^{1/t}-1\times 100\%$ (11)
During the training period, the most profitable parameter set is〈open threshold = 1.8 $z$-score, close threshold = 0.4 $z$-score, window size = 900 intervals⟩. Some example results of the grid search are presented in Table 3.
OPEN_THRES CLOS_THRES PERIOD RTOT (%)
4.0 2.0 2000 0.0651
4.0 0.5 500 0.5024
3.0 1.0 500 0.9993
3.0 0.5 1000 0.8932
3.0 0.5 500 1.0704
2.5 0.3 700 2.1542
2.5 0.5 700 1.5667
3.0 0.3 700 1.3160
2.1 0.4 700 2.5633
2.1 0.3 800 2.6916
2.3 0.4 800 2.3096
2.1 0.4 800 2.8202
2.0 0.4 1000 2.7339
2.0 0.4 900 3.0400
1.9 0.3 900 2.8989
1.9 0.4 900 3.1077
1.8 0.4 900 3.0565
Table 3: Trading Parameters Tuning
The setup of RL-based pair trading relies on these parameters. The window size decides the retrospective length of the spread, and the thresholds decide the zones. Algorithms such as PPO and A2C are applicable to both discrete and continuous action spaces. Some algorithms, e.g. DQN, can only be used on discrete space, and DDPG is only applicable in a continuous space. Therefore, we adopt PPO, DQN and A2C on RL pair trading that decides the timing, and PPO, A2C,SAC on RL pair trading that decides both timing and investment quantity. The algorithms are adopted from the Baseline3 collection [Raffin et al., 2021].
5.1.3 Evaluation Metrics
Our main concern is the highest profitability in trading techniques. We care about the cumulative return, which is the total profit for the testing period, as well as the annualised return Compound Annual Growth Rate (CAGR). With $V(t_{0})$ as the initial state, $V(t_{n})$ as the final state, $t_{n}-t_{0}$ is period of trading in years, the CAGR is (Equation 12):
$\operatorname{CAGR}\left(t_{0},t_{n}\right)=\left(\frac{V\left(t_{n}\right)}{V\left(t_{0}\right)}\right)^{\frac{1}{t_{n}-t_{0}}}-1$ (12)
There are some popular indicators for distinguishing whether a strategy is profit-risk effective, eg. Sharpe Ratio. Sharpe Ratio [Sharpe, 1964] where $R_{p}$ is the return of the trading strategy, $R_{f}$ is the interest rate [^4] We adopt the Federal Reserve interest rate 5.5% as on 13 June 2024, and $\sigma_{p}$ is the standard deviation of the portfolio’s excess return (Equation 13):
Sharpe $\text{Sharpe Ratio}=\frac{R_{p}-R_{f}}{\sigma_{p}}$ (13)
We also care about the strategies’ activities, such as order count and win/loss ratio.
The indicators used for comparison are presented in Table 4.
Profitability Indicator Description
Cumulative Return Profit achieved during trading period
CAGR Compound Annual Growth Rate
Sharpe Ratio Risk-adjusted returns ratio
Activity Indicator Description
Total Action Count Total orders executed
Win/Loss Action Count Number of winning/losing trades
Win/Loss Action Ratio Ratio of winning to losing trades
Max Win/Loss Action Maximum profit/loss per Action in Bitcoin
Avg Action P&L Average profit/loss per trade in Bitcoin
Time in Market Percentage of time invested in the market
Risk Indicator Description
Volatility (ann.) Annualized standard deviation of returns
Skew Asymmetry of returns distribution
Kurtosis “Tailedness” of returns distribution
Table 4: Descriptive Table of Evaluation Metrics
5.2 Experimental Results
We present the profitability and risk results from our experiments along with the trading indicators.
5.2.1 Result Comparison
Our work is compared with standard pair trading (Section 2.1) and state-of-the-art pair trading techniques (Section 3.2).
Table 5: Evaluation Metrics Comparison between Trading Techniques
Gatev et al. (2006) Kim & Kim (2019) RL1 RL2
RL algo. NA PPO A2C DQN PPO A2C DQN PPO A2C SAC
Profitability
Cumulative Return 8.33% -0.16% -35.16% -35.79% 1.89% 9.94% -31.99% -77.81% 31.53% -87.12%
CAGR 195.12% -2.19% -99.71% -99.75% 30.05% 278.72% -99.56% -100.00% 3974.65% -100.00%
Sharpe Ratio 25.91 -1.67 -2.04 -2.60 5.44 32.74 -8.77 -1.99 94.34 -1.93
Activities
Total Action Count 490 43 1248 1062 1304 249 879 3443 229 2798
Won Action Count 284 24 600 503 578 240 232 842 162 917
Lost Action Count 206 19 648 559 726 9 647 2601 67 1881
Win/Loss Action Ratio 1.38 1.26 0.93 0.90 0.80 26.67 0.36 0.32 2.42 0.49
Max Win Action (\$) 75.35 163.52 606.75 606.75 43.72 121.74 70.59 307.78 648.87 160.15
Max Loss Action (\$) -27.73 -187.86 -763.70 -553.25 -108.51 -21.33 -282.84 -389.22 -64.97 -1456.43
Avg Win Action P&L (\$) 14.50 41.72 38.68 37.47 5.51 17.77 11.30 15.88 90.94 8.03
Avg Loss Action P&L (\$) -2.90 -54.68 -58.75 -60.78 -3.28 -7.06 -24.95 -17.78 -21.00 -23.49
Risk
Volatility (ann.) 6.01% 3.93% 51.43% 40.43% 3.61% 6.30% 11.92% 53.04% 27.30% 54.66%
Skew 1840 -54 -358 -358 -874 2673 -1899 -374 4314 -3048
Kurtosis 48145 135 4138 4201 133603 114944 54808 12987 254283 138851
  • 1
    The trading period is from 01-Dec-2023 to 31-Dec-2023.
  • 2
    The transaction cost is 0.02%, and the interest rate is 5.5%.
  • 3
    RL 1 stands for the pair trading that allows Reinforcement Learning to decide upon the investment timing.
  • 4
    RL 2 stands for Reinforcement Learning pair trading that allows the RL agent to decide both investment timing and quantity.
(a) Gatev et al. (2006) Pair Trading
(b) Kim & Kim (2019) Pair Trading (PPO)
(c) RL 1 Pair Trading (A2C)
(d) RL 2 Pair Trading (A2C)
Figure 6: The portfolio value trend comparison between agents with the best performance
Our results are presented in Table 5. The results display a positive return for traditional pair trading technique Gatev et al. [2006]. The algorithm A2C displays a positive return for RL pair trading techniques. However, algorithms PPO, SAC, and DQN do not perform as well as A2C. If we view A2C as the chosen algorithm for pair trading, the results show a steady income from pair trading. The traditional pair trading of Gatev et.al. 2006 displays a stable income compared to others due to its rule-based execution stability.
The first adoption of RL 1 pair trading is close to the traditional method. The result table shows that it achieved much better results than the traditional pair trading approach Gatev et al. [2006]. The second adoption of RL 2 pair trading is significantly different from the RL 1 trading that decides only timing, which produces more profit than other techniques under the same level of volatility. Kim & Kim (2019)’s method did not achieve a positive return. Since the method was developed for the forex market, it has not adapted well to the extremely volatile cryptocurrency world.
Behaviour-wise, PPO, DQN and SAC tend to conduct excessive transactions that are not profitable. On the contrary, A2C have fewer trades but higher profits on each trade. The RL 2 pair trading shows further fewer total actions because of the adjust position action, where we do not consider a position adjustment as one trade until it is closed. Apart from the result in Table 5, the portfolio growth trend with the best performing RL algorithm agent is presented in Figure 6. Most of the pair trading experiments, including Gatev et al. [2006], RL1, and RL2, display a stable upturn, which is ideal from the perspective of pair trading. We can see from the drawdown graphs that RL 1 produces fewer drawdowns compared to the non-RL pair trading method from Gatev et al. [2006], with even more profit. RL2 displays the strongest profitability as well as the largest drawdown. In general, all three pair trading methods show the ability to generate stable income in a volatile trading market.
5.2.2 Effect of Transaction Cost
The profitability of high-frequency trading techniques always receives a significant impact from transaction costs. Cryptocurrency exchanges normally provide a large-volume discount scheme. The Binance fee ranges from 0.02% to even 0%, depending on the volume and the holding of their membership token. Considering that the users of these techniques may benefit from different transaction fee tiers, we explore the trading techniques under different transaction cost tiers as well.
Indicators Trading approaches
0.05% Transaction Fee Gatev et al. (2006) Kim & Kim (2019) RL1 RL2
Cumulative Profit 5.02% -0.26% 5.76% 7.40%
Sharpe Ratio 14.60 -2.34 21.00 7.82
Total Action Count 490 43 154 207
Won Action Count 246 23 152 110
Lost Action Count 244 20 2 97
Win/Loss Action Ratio 1.01 1.15 76.00 1.13
Max Win Action (\$) 72.82 114.76 43.36 606.22
Max Loss Action (\$) -30.26 -169.52 -8.30 -168.81
Avg Win Action P&L (\$) 13.57 37.94 16.10 70.99
Avg Loss Action P&L (\$) -4.99 -47.68 -5.64 -48.28
0.01% Transaction Fee Gatev et al. (2006) Kim & Kim (2019) RL1 RL2
Cumulative Profit 9.43% -1.13% 9.88% 33.99%
Sharpe Ratio 29.84 -7.07 33.24 104.40
Total Action Count 490 43 251 181
Won Action Count 317 20 242 149
Lost Action Count 173 23 9 32
Win/Loss Action Ratio 1.83 0.87 26.89 4.66
Max Win Action (\$) 76.20 65.99 121.74 675.23
Max Loss Action (\$) -26.88 -169.66 -21.33 -27.91
Avg Win Action P&L (\$) 13.93 24.88 17.52 98.74
Avg Loss Action P&L (\$) -2.48 -44.81 -7.06 -10.79
0% Transaction Fee Gatev et al. (2006) Kim & Kim (2019) RL1 RL2
Cumulative Profit 10.54% -2.00% 9.94% 80.92%
Sharpe Ratio 33.90 -5.76 32.74 2668.86
Total Action Count 483 43 249 429
Won Action Count 363 23 240 342
Lost Action Count 120 20 9 87
Win/Loss Action Ratio 3.02 1.15 26.67 3.93
Max Win Action (\$) 77.04 163.59 121.74 699.51
Max Loss Action (\$) -26.03 -217.54 -21.33 -72.21
Avg Win Action P&L (\$) 13.07 36.69 17.77 104.25
Avg Loss Action P&L (\$) -2.43 -80.76 -7.06 -16.68
Table 6: Evaluation Metrics Comparison under Different Transaction Cost
  • 1
    The trading period is from 01-Dec-2023 to 31-Dec-2023 with interest rate as 5.5%.
  • 2
    RL 1 stands for the pair trading that allows Reinforcement Learning to decide upon the investment timing.
  • 3
    RL 2 stands for Dynamic Scaling Reinforcement Learning pair trading that allows the RL agent to decide both investment timing and quantity.
  • 4
    We adopted algorithm PPO for Kim & Kim (2019), and A2C for RL1 and RL2.
With more exploration under 0.05%, 0.01% and 0% transaction costs compared to the default 0.02% transaction cost, we can see the significant impact of decreasing the transaction cost. The participating approaches are adopted with the most profitable algorithm based on the backtesting result (Table 5). We can see that trading techniques generally perform better under lower transaction costs. RL-based techniques tend to perform more trades when the transaction costs are lower.
6 Discussion and conclusions
Pair trading has been a popular algorithmic trading method for decades. The in-demand high-frequency trading domain requires a fast-track decision-making process. However, the traditional rule-based pair trading technique lacks the flexibility to cater to volatile market movement. In this research, we proposed a mechanism to adopt Reinforcement Learning (RL)to observe the market and produce profitable pair trading decisions. The the first adoption of Reinforcement Learning pair trading grants the RL 1 agent the flexibility to decide action timing. The second adoption of Reinforcement Learning 2 Pair Trading further gives RL agent the access to decide the timing and invest quantity.
We compared it to the traditional rule-based Pair Trading technique [Gatev et al., 2006] and a state-of-the-art RL pair trading technique [Kim & Kim, 2019], for 2023-Dec in the cryptocurrency market for BTCEUR and BTCGBP under standard future 0.02% transaction cost. Kim & Kim’s method does not perform well in the cryptocurrency world. The Gatev et al.’s method achieved 8.33% per trading period. our first adoption of RL 1 method achieved 9.94%, and the second adoption of RL 2 method achieved 31.53% returns during the trading period. The outperformance is generally consistent across different transaction costs. The evaluation metrics show that RL-based techniques are generally more active than traditional techniques in the cryptocurrency market under various transaction costs. In general, our trading methods have greater market participation than Gatev et al.’s traditional rule-based pair trading and Kim & Kim’s threshold-adaptive RL Pair Trading (Table 56).
Comparison between RL-based pair trading revealed the relationship between profitability and actions. Because financial trading is a special case of RL environment, every action in financial trading is punished by the transaction cost. We notice that profitable RL trading often has less total trade count and higher profit per win trade. That means the RL is better at spotting chances to make higher profits. The RL 2 pair trading produces higher profits because of higher average wins from the position adjustment mechanism.
The techniques open some future work opportunities. We can develop the RL pair trading with applications to multi-leg strategies. At this moment, pair trading either relies on forming a pool of assets into a two-leg pair and executing trades upon the selected pair or relies on an optimisation-based analytical function for pair combining them amongst a selected pool. We believe the RL optimiser might provide better execution of pair formation. The breakdown of Pair Formation from preliminary preparation into part of trading could significantly boost the trading efficiency, diversify the asset holding risk and potentially enhance profitability.
Acknowledgements