### v2

# pair trading; reinforcement learning; algorithmic trading; deep learning; cryptocurrency

Abstract
Cryptocurrency is a cryptography-based digital asset with extremely volatile prices. Around USD 70 billion worth of cryptocurrency is traded daily on exchanges. Trading cryptocurrency is difficult due to the inherent volatility of the crypto market. This study investigates whether Reinforcement Learning (RL) can enhance decision-making in cryptocurrency algorithmic trading compared to traditional methods. In order to address this question, we combined reinforcement learning with a statistical arbitrage trading technique, pair trading, which exploits the price difference between statistically correlated assets. We constructed RL environments and trained RL agents to determine when and how to trade pairs of cryptocurrencies. We developed new reward shaping and observation/action spaces for reinforcement learning. We performed experiments with the developed reinforcement learner on pairs of BTC-GBP and BTC-EUR data separated by 1 min intervals ($n$ = 263,520). The traditional non-RL pair trading technique achieved an annualized profit of 8.33%, while the proposed RL-based pair trading technique achieved annualized profits from 9.94% to 31.53%, depending upon the RL learner. Our results show that RL can significantly outperform manual and traditional pair trading techniques when applied to volatile markets such as cryptocurrencies.
keywords:
pair trading; reinforcement learning; algorithmic trading; deep learning; cryptocurrency
[…]
1 Introduction
Arbitrage is a subdomain of financial trading that profits from price discrepancies in different markets (Dybvig and Ross, 1989). Pair trading is one of the well-known arbitrage trading methods in financial markets. Arbitrageurs identify two highly correlated assets to form a pair. When a price discrepancy happens, they buy the underpriced asset and sell the overpriced correlated asset to profit from the mean reversion of the prices. With the rise of high-frequency trading, the ability to conduct fast and accurate analyses has become critical. Arbitrage requires practitioners to constantly analyze the market conditions at the fastest speed possible, as arbitrageurs must compete for transitory opportunities (Brogaard et al., 2014). Therefore, we explore how Artificial Intelligence (AI)can enhance the process of pair trading, focusing on the speed and adaptability of decision-making.
Reinforcement Learning (RL)is a captivating domain of AI. The idea of RL is to let the agent(s) learn to interact with an environment. The agent should learn from the environment’s responses to optimize its behavior (Sutton and Barto, 2018). If we view the financial market from the perspective of the RL environment, actions in the financial market are investment decisions. By allowing agents to adapt dynamically to market conditions, RL has the potential to overcome the limitations of static, rule-based strategies in volatile and complex financial environments. For gaining profits, arbitrageurs are incentivized to train agents to produce lucrative investment decisions, and RL facilitates agents’ learning process from the profit/loss of the market.
The combination of RL and various financial trading techniques is still evolving rapidly. There has been some work in RL infrastructural construction (Liu et al. 2021, 2022a, 2022b) and some experiments in profitable RL agent training (Meng and Khushi, 2019; Zhang et al., 2020; Pricope, 2021). Trading actions in traditional pair trading follow static rules. In reality, the complexity of financial markets should allow more flexibility in the decision-making process. An experienced trader might analyze market conditions to make informed decisions. However, it is not feasible to output efficient decisions at short, intermittent intervals 24/7. RL algorithms enable a fast-track decision-making process for analyzing trading signals and generating trading actions.
Designing a high-frequency trading system based on RL requires addressing critical challenges. The first challenge is how to construct an RL environment that accommodates RL agents for arbitrage. The second challenge involves identifying compatible instruments with historical correlations to form profitable pairs. The third challenge concerns timing. Instead of blindly following preset rules, the system requires flexibility in choosing investment timings to achieve greater profitability. The final challenge involves investment quantity. Since investment opportunities vary in quality, a critical consideration is whether RL agents can replicate decision-making capabilities comparable to the scrutiny applied by experienced traders.
This paper investigates key questions centered around the application of Reinforcement Learning (RL) in pair trading. To address the fast decision-making requirements in a high-frequency trading environment, we constructed a tailored RL environment for pair trading and fine-tuned reward shaping to encourage the agent to make profitable decisions. The contributions of this work are as follows: (1) the construction of an RL environment specifically designed for quantity-varying pair trading; (2) the proposal of a novel pair trading method that incorporates adaptive investment quantities to capture opportunities in highly volatile markets; (3) the use of a grid search technique to fine-tune hyperparameters for enhanced profitability; (4) the introduction of an RL component for market analysis and decision-making in pair trading, along with a novel RL model optimized for investment quantity decisions.
The structure of the paper is arranged as follows: the background and related work are introduced in Sections 2 and 3. The methodology is presented in Section 4. Experiments and results are included in Section 5. A discussion of the results and conclusions is provided in Section 6.
2 Background
First, we define the basic terms of financial trading. A long position is created when an investor uses cash to buy an asset, and a short position is created when an investor sells a borrowed asset. The portfolio is the investor’s total holding, including long/short position and cash. Transaction cost is a percentage fee payable to the broker for any long/short actions. Finally, risk is defined as the volatility of the portfolio.
2.1 Traditional Pair Trading
Classical pair trading consists of two distinct components known as legs . A leg represents one side of a trade in a multi-contract trading strategy. Under the definition of pair trading, “longing the first asset and shorting the second asset” is called a long leg , and “shorting the first asset and longing the second asset” is called a short leg . The two assets are always bought and sold in opposite directions in pair trading. Therefore, the overall pair trading strategy is considered to be market neutral because the profits from the long position and the short are offset by the direction of the overall market. Gatev et al.’s (2006) work is the most cited traditional pair trading method. It follows the OODA (observe, orient, decide, and act) loop (Fadok et al., 1995). Before entering the market, the first step is to choose the proper assets in a pair. The Sum of Squared Deviation (SSD)is the measurement calculated from the prices of assets $i$ and $j$. Through exhaustive searching in a formation period $T$, the assets with the smallest SSD are bound as a pair Equation (1).
$SSD_{p_{i},p_{j}}=\sum_{t=1}^{T}(p_{i}-p_{j})^{2}.$ (1)
  • Observe is the process of market analysis. The price of assets in pairs is collected and processed. The price difference $(p_{i}-p_{j})$ is called spread $S$. The arbitrageurs observe the current positions and spread of the current market.
  • Orient is the process exploring what could be done. Three possible actions for pair trading are long leg, short leg, and close position, as defined above.
  • Decide what action to take. Position opening triggers when the price difference deviates too much. This is indicated by the spread movement beyond an open threshold. Position closing happens when the spread reverts back to some closing threshold. Gatev et al. (2006) adopted two times the standard deviation of the spread as the opening threshold and the price crossing as the closing threshold. In practice, the threshold varies according to the characteristics of the financial instrument.
  • Act once the decision is made. The long leg orders us to buy asset $i$ and sell asset $j$. The short leg orders us to sell asset $i$ and buy asset $j$. Closing a position means clearing all the active positions to hold cash only.
A graphical visualization of pair trading is presented in Figure 1. Figure 1a shows the market interactions according to the Spread (S) and thresholds. A position is opened whenever the spread deviates beyond the open threshold. The position closure happens when the spread reverts below the close threshold. Figure 1b, which shares the same time axis with (a), is a stretched view of (a). It presents the corresponding actions with the crossing of Spread (S) and zones. The spread deviations are classified into zones based on the Spread ($S$), Open-Threshold ($OT$), and Close-Threshold ($CT$) Equation (2):
Figure 1: Stretched pair trading view of price distance between $p_{i}$ and $p_{j}$. Figure (b), which shares the same time axis with (a), is a stretched view of (a). It presents the corresponding same actions with the crossing of Spread (S) and zones in two different views.
$\left\{\begin{array}[]{rl}\text{Short Zone:}&+OT<S<+\infty,\\ &\text{spread deviates beyond open threshold}\\ \text{Neutral Short Zone:}&+CT<S<+OT,\\ &\text{spread deviates between open and close threshold}\\ \text{Close Zone:}&-CT<S<+CT,\\ &\text{spread reverts between close thresholds}\\ \text{Neutral Long Zone:}&-OT<S<-CT,\\ &\text{spread deviates between open and close threshold}\\ \text{Long Zone:}&-\infty<S<-OT,\\ &\text{spread deviates below open threshold}\\ \end{array}.\right.$ (2)
2.2 Reinforcement Learning
Reinforcement Learning (RL)is used to train an agent to maximize rewards while interacting with an environment (Sutton and Barto, 2018). The environment for RL is required to be a Markov Decision Process (MDP(Bellman, 1957), which means it is modeled as a decision-making process with the following elements〈State ($S$), Action ($A$), Transition ($P_{A}$), Reward ($R_{A}$)⟩. The goal is to train the agent to develop a policy ($\pi$) that fulfills an objective, e.g., maximizing profits in a trade. At every trading interval $t$, according to state $S$, that the agent observes, action $A$ is chosen based on policy $\pi$. The environment rewards/punishes the state transition of $S_{t}\rightarrow S_{t+1}$ with environment reward $r$. If we assume $\gamma$ to be the discount factor for the time-value discount of future reward, RL trains a policy $\pi$ that maximizes the total discounted reward $G_{t}$, as shown in Equation (3):
$G_{t}=\sum_{i=0}^{\infty}\gamma^{i}r_{t+i}.$ (3)
RL algorithms can broadly be classified based on three criteria: value/policy-based, on/off policy, and actor/critic network (AlMahamid and Grolinger, 2021). Value-based methods estimate state-action value functions for decision-making. Policy-based methods directly learn action selection policies. The on-policy method requires data generated by the current policy, and the off-policy method is capable of leveraging past experiences from potentially different policies. Moreover, actor–critic architectures, where an actor–network proposes actions and a critic network evaluates, have shown a better performance in facilitating policy improvement through this feedback loop. Most recent research favors an actor–critic architecture instead of actor-only or critic-only methods for better performance (Meng and Khushi, 2019; Zhang et al., 2020). Therefore, only actor–critic algorithms are adopted in this study.
Based on the RL classification criteria, some representative algorithms have been selected for this study including Deep Q-Learning (DQN)(Mnih et al., 2013), Soft Actor Critic (SAC)(Haarnoja et al., 2019), Advantage Actor-Critic (A2C)(Sutton and Barto, 2018), Proximal Policy Optimization (PPO)(Schulman et al., 2017). Diversified RL algorithms are experimented with to choose the most effective one in pair trading.
3 Related Work
3.1 Reinforcement Learning in Algorithmic Trading
Reinforcement learning in AlphaGo captured the world’s attention in 2016 by participating in a series of machine versus human competitions on the board game GO (Silver and Hassabis, 2016). Surprisingly, the research regarding RL in the financial market started long before that. Recurrent reinforcement learning studies were the mainstream works (Gold, 2003; Bertoluzzo and Corazza, 2007; Maringer and Ramtohul, 2012; Zhang and Maringer, 2016) in the early stage of financial trading. After the upsurge of AlphaGo, some significant advancements were brought to RL trading as well; Huang (2018) re-described the Markov Decision Process (MDP)financial market as a game process to incorporate RL as a financial trading game (Huang, 2018). Pricope (2021) proposed deep RL agents to develop profitable high-frequency trading strategies with sequential model-based optimization tuning the hyperparameters. With the recent development, newer RL models such as Deep Q-Learning (DQN), Policy Gradients (PG), and Advantage Actor-Critic (A2C)have also been introduced by researchers (Meng and Khushi, 2019; Zhang et al., 2020; Mohammadshafie et al., 2024) for financial trading. A noteworthy research work is that of the FinRL group in the infrastructures and ensemble learning mechanism (Liu et al. 2021, 2022a, 2022b).
3.2 Reinforcement Learning in Pair Trading
Reinforcement Learning, in combination with pair trading, is not an untapped domain. RL has ameliorated multifaceted aspects of the traditional method of pair trading brought up by Gatev et al. (2006). The RL technique Ordering Points To Identify The Clustering Structure contributed to the pair selection stage by leveraging a clustering algorithm to produce better pair choices (Sarmento and Horta, 2020). Vergara and Kristjanpoller (2024) traded deep reinforcement learning in the cryptocurrency market with ensemble practice, combining classical pair trading with the RL framework. reCurrent Reinforcement lEarning methoD for paIrs Trading (CREDIT), an algorithm that takes into consideration both profitability and risks, was engineered by Han et al. (2023). Reward shaping is also an interesting area where some work has been done in RL trading (Lucarelli and Borrotti, 2019; Wang et al., 2021). Kim and Kim’s work in (2019) is the most recent RL pair trading method. Their focus is on utilizing RL to find the most trading opportunities. Instead of fixed thresholds, the RL agent in Kim and Kim’s work produces thresholds for the upcoming trading period. Open, close, and stop-loss thresholds determine the profits of pair trading.
Our work introduces a novel method to combine RL with pair trading. The work of Gatev et al. (2006) is not efficient enough for a high-frequency market. The state-of-the-art method of Kim and Kim (2019) has some deficiencies: (i) it requires the market’s volatility to be relatively stable. The RL agent may produce unsuitable thresholds if the market experiences increased volatility. (ii) It lacks flexibility in the investment amount. Opportunities with different qualities are programmed to be invested with the same amount of capital. Once the RL agent determines a threshold, the trading algorithm executes a trade at pre-determined thresholds. We leverage RL to make investment timing and quantity decisions. The adjustable investment amount is a novel feature of our RL pair trading. An RL agent measures how well the investment opportunities are based on observations and invests a larger amount in more promising market conditions. Having another dimension on the investment side should further enhance profitability and reduce risks.
4 Methodology
In this section, we introduce the architecture of the methodology (Figure 2). The architecture includes five steps: (1) pair formation for selecting assets to form a tradeable pair (Section 4.1); (2) spread calculation utilizing the moving-window technique to extract the spread in a limited retrospective time frame (Section 4.2); (3) parameter selection from an historical dataset to decide the most suitable hyperparameters for pair trading (Section 4.2); (4) RL trading by allowing RL to decide the trading timing and quantity in pair trading (Section 4.4); (5) investment action for taking the actions produced from RL trading into market execution.
Figure 2: Architecture of trading strategies.
4.1 Pair Formation
Pairs are selected based on two criteria: correlation and cointegration. The widely adopted Pearson’s correlation (Perlin, 2007; Do and Faff, 2010) is given by
$\rho_{X,Y}=\frac{\mathrm{cov}(X,Y)}{\sigma_{X}\sigma_{Y}},$ (4)
where $\rho_{X,Y}$ is the correlation coefficient between assets $X$ and $Y$, $\mathrm{cov}(X,Y)$ is the covariance of $X$ and $Y$, and $\sigma_{X}$ and $\sigma_{Y}$ are the standard deviations of $X$ and $Y$, respectively. The Engle–Granger cointegration test (Burgess, 2003; Dunis and Ho, 2005) involves two steps. First, the linear regression is performed:
$Y_{t}=\alpha+\beta X_{t}+\epsilon_{t},$ (5)
where $Y_{t}$ and $X_{t}$ are the asset price series, $\alpha$ and $\beta$ are the regression coefficients, and $\epsilon_{t}$ is the residual term. The second step tests the residuals $\epsilon_{t}$ for stationarity using an Augmented Dickey-Fuller (ADF)(Dickey and Fuller, 1979) test. The ADF test regression is given in Equation (6):
$\Delta\epsilon_{t}=\gamma\epsilon_{t-1}+\sum_{i=1}^{p}\delta_{i}\Delta\epsilon_{t-i}+\nu_{t},$ (6)
where $\Delta\epsilon_{t}$ is the first difference of the residuals, $\gamma$ is the coefficient to be tested for stationarity, $p$ is the number of lagged difference terms included, and $\nu_{t}$ is the error term. If $\gamma$ is significantly different from zero, the residuals are stationary, indicating co-integration.
A moving window is applied to historical pricing data, as shown in Figure 3. In this figure, the blue line represents the historical prices, while the dashed boxes illustrate the moving window. During the selection phase, averaged correlation and co-integration batches are employed to ensure that the selected assets exhibit a strong, long-term statistical relationship.
Figure 3: Window-size cut for correlation and co-integration testing.
4.2 Spread Calculation
The second step is a moving window mechanism to capture the spread movement (Figure 2 Step-2). Spread $\epsilon_{t}$ is calculated at every selected trading interval (e.g., every five minutes) and is the error term $s$ from a regression between the two prices $p_{i}$ and $p_{j}$. $\beta_{0}$ and $\beta_{1}$ are the intercept and slope, respectively, which follow a normal distribution with a mean of 0 and a standard deviation of $\sigma$ Equation (7):
$p_{i}=\beta_{0}+\beta_{1}\times p_{j}+s_{i}\sim N(0,\sigma^{2}).$ (7)
We normalize the spread with $z$-score Equation (8) to scale the spread into constant mean and standard deviation. The mean of the spread in the sliding window is represented as $\bar{s}$:
$Z=\frac{{s-\bar{s}}}{{\sigma_{s}}}.$ (8)
4.3 Parameter Selection
Three parameters to be explored are window size, open threshold, and close threshold. Window size $\in\mathbb{Z}^{+}$ is the number of historical samples in the moving window. Thresholds $\in\mathbb{Q}^{+}$ are the entry and exit signals of trading actions that are highly linked to market conditions.
Excessively wide thresholds suit more volatile markets, and conservatively narrow thresholds result in smaller but steadier wins. The combination of parameters of the highest profitability 〈Window Size , Open Threshold , Close Threshold ⟩ are selected from a search pool through a grid search in practice. Gatev et al. (2006) adopted 2 times the standard deviation as the open threshold and the deviation crossing point as the close threshold (Figure 1). However, on the one hand, the parameters should vary with the arbitrage instruments as well as the market condition. Hence, the window-sliding mechanism is incorporated to reflect the heterogeneity of the pricing variance (Mandelbrot, 1967).
4.4 Reinforcement Learning Pair Trading
After we run the grid search of window-sliding pair trading, the next problem concerns “when” and “how much” to trade. Pair trading results from following pre-set rules (à la Gatev et al. 2006), which are obtainable using window-sliding pair trading. However, we want to know if RL produces better investment decisions than blindly following the rules. Therefore, the most profitable parameter combination is passed onto further RL-based pair trading so that we can compare the results between RL-based pair trading and non-RL pair trading.
4.4.1 Observation Space
Observation space stands for the information an RL agent observes. The agent observes market information to make decisions. The observations adopted for our RL environment are as follows: 〈Position, Spread, Zone⟩.
  • Position $\in[-1,1]$: Position stands for the current portfolio value. Position is a percentage measuring the direction of investment (c.f. Figure 4). As in pair trading, we define longing the first asset with shorting the second asset as ‘holding a long leg’ and the other way around as ‘holding a short leg’. Assuming that we do not use leverage, holding a long leg with a 70% portfolio value gives Position = 0.7. Holding a short leg with a 30% portfolio value gives Position = --0.3. Position 0 means we only hold cash.
  • Spread $\in\mathbb{R}$: This represents how much the current spread has deviated from the mean (Section 4.2).
  • Zone $\in\{\text{Zones}\}$: Zone is an important indicator that comes from the comparison between the $z$-score with the thresholds for signals (Figure 1b). Traditional pair trading (Gatev et al., 2006; Yang and Malik, 2024) takes the zone as the direct trading signal. However, in RL-based pair trading, the zone is an observation for the RL agent to make better decisions.
Figure 4: The value of position observation based on investment.
4.4.2 Action Space
Since the pair trading technique is a relatively low-risk strategy, most applications invest with a fixed amount or the complete portfolio value (Burgess, 2003; Perlin, 2009; Huck, 2010). Meanwhile, it is natural for an experienced trader to invest different amounts based on the quality of opportunities. Opportunities with a higher probability of success are worth more investment. Therefore, we investigate granting the RL agent not only the decision of when to invest but also the freedom to choose the investment amount.
Take $A\in[-1,1]$ as the action. Similar to the observation space (c.f. Figure 4), the action ranges from --1 to 1, representing the investment as a percentage of the portfolio value in the long leg and short leg directions. Investing 50% of the portfolio as a long leg means A = 0.5. Investing 20% of the portfolio in the short leg means A= --0.2.
In practice, we have to consider the relationship between the existing position and the next action. We classify the execution of action as open position , adjust position , or close position .
  • Open position is the action of opening a new position.
  • Close position is the closure of a position.
  • Adjust position happens when a previous position is open, and the RL agent wants to open another position. For example, if the current position is a 70% long-leg and the new action is A = 0.8, only the extra 10% shall be actioned.
4.4.3 Reward Shaping
The RL reward consists of three components: action reward , portfolio reward , and transaction punishment .
  • Portfolio reward is the profit/loss from closing a position. The portfolio value $V_{p}$ only updates when the position closes. $V^{\prime}_{p}$ is the position value at the start of a trading period $p$. Then, upon closing the trade at the end of trading period $p$, the reward it calculated, as shown in Equation (9).
    Profit $\text{Profit Reward}=V_{p}-V^{\prime}_{p}.$ (9)
  • Action reward means the agent needs to be rewarded for taking a desired action in the corresponding zone. In general, the agent is free to decide on any action. However, we use action reward to encourage the agent to choose desired actions. It rewards the agent for making a desired action in certain zones (Table 1) with some freedom in neutral zones. The stronger the action reward, the more it resembles traditional pair trading.
  • Transaction punishment is a negative reward for encouraging small adjustments instead of large changes in the position. The punishment is the difference between the action and position. If the current position in observation is $P$ and the action is $A$, the transaction punishment is Equation (10):
    Transaction $\text{Transaction Punishment}=P-A.$ (10)
Table 1: Rewarding behaviors in zones.
                  Zones                   Rewarding Behavior
                  Short Zone                   Short leg
                  Neutral Short Zone                   Short leg or Close
                  Close Zone                   Close
                  Neutral Long Zone                   Long leg or Close
                  Long Zone                   Long leg
4.5 Dynamic Agents
Trading agents in RL environments operate by taking actions within a defined action space based on observations from the state space. This section details the design of two dynamic agent settings, RL1 and RL2, each tailored to address specific aspects of pair trading. These agents leverage the Markov Decision Process (MDP) framework to model the complexities of financial markets, enabling adaptive decision-making in volatile conditions.
In the first setting, RL1 is tasked with optimizing trade timing and directionality. Pair trading is modeled as a Markov Decision Process (MDP), represented as $(\mathcal{S}_{1},\mathcal{A}_{1},\mathcal{T}_{1},r_{1},\gamma_{1})$:
  • $\mathcal{S}_{1}$ represents the state space, including normalized price spreads, historical z-scores, and zone indicator.
  • $\mathcal{A}_{1}$ defines a discrete action space consisting of three possible actions: initiating a long–short position, closing existing positions, or initiating a short–long position. This allows the agent to determine the optimal direction and timing for trades.
  • $r_{1}(s,a)$ is the reward function, defined as follows:
    $r_{1}(s_{t},a_{t})=\begin{cases}\Delta P_{t}-c,&\text{if a trade is executed;}\\ 0,&\text{if no trade is executed,}\end{cases}$
    where $\Delta P_{t}$ represents the profit or loss from the trade, and $c$ is the transaction cost. The reward function penalizes the agent for transaction costs while directly linking rewards to trade profitability.
  • The goal is to maximize the cumulative discounted reward:
    $R_{1}=\mathbb{E}\left[\sum_{t=0}^{\infty}\gamma_{1}^{t}r_{1}(s_{t},a_{t})\right].$
RL2 extends RL1 by shifting focus from trade timing to determining the investment quantity for a given trade opportunity. It models pair trading as an MDP defined by $(\mathcal{S}_{2},\mathcal{A}_{2},\mathcal{T}_{2},r_{2},\gamma_{2})$:
  • $\mathcal{S}_{2}$ is the state space, which is identical to RL1.
  • $\mathcal{A}_{2}=[-1,1]$, where the continuous value represents the investment quantity. Here, 00 stands for no involvement, positive values represent buying, and negative values represent selling.
  • $r_{2}(s,a)$ is the reward function, defined as follows:
    $r_{2}(s_{t},a_{t})=\Delta P_{t}\cdot a_{t}-c(|a_{t}|),$
    where $c(|a_{t}|)$ represents transaction costs proportional to the absolute investment size $|a_{t}|$. This reward structure incentivizes the agent to optimize both the direction and magnitude of its investment.
  • The objective is to maximize the cumulative discounted reward:
    $R_{2}=\mathbb{E}\left[\sum_{t=0}^{\infty}\gamma_{2}^{t}r_{2}(s_{t},a_{t})\right].$
The primary differences between RL1 and RL2 lie in their action spaces and reward functions. RL1 operates with a discrete action space and focuses on optimizing directional timing and trade management. In contrast, RL2 uses a continuous action space $\mathcal{A}_{2}=[-1,1]$, enabling it to adjust investment sizes dynamically. The environments are designed to guide the agents by rewarding profitable actions and penalizing costly ones, encouraging effective decision-making for timing and quantity. The exact mechanisms driving these decisions are embedded within the neural network, shaped by the agent’s interactions with the environment.
5 Benchmark Results
Next, we carry out experiments using the proposed methodology. We adopt the same dataset and the same parameters for non-RL pair trading and RL pair trading for comparison purposes.
5.1 Experimental Setup
We experiment with window-sliding pair trading and RL pair trading in the cryptocurrency market. The cryptocurrency market is famous for its volatility, easy access, and 24/7 operating time. Data preprocessing follows two steps. Firstly, the evaluation of the correlations among trading pairs for selection of arbitrage candidates; next, searching through the combination of thresholds and retrospective periods for suitable signals.
5.1.1 Datasets
[…]
Figure 5: Prices of BTCEUR and BTCGBP.
Table 2: Correlation and co-integration of pair formation.
Pairs 1m 3m 5m
coint corr coint corr coint corr
BTCEUR-BTCGBP 0.5667 0.8758 0.4667 0.8759 0.4667 0.8754
BTCEUR-BTCRUB 0.3333 0.8417 0.3333 0.8417 0.3167 0.8416
BTCEUR-BTCUSD 0.1667 0.9328 0.2000 0.9327 0.2000 0.9329
BTCGBP-BTCRUB 0.3500 0.7606 0.3333 0.7608 0.3333 0.7603
BTCGBP-BTCUSD 0.4833 0.8404 0.4167 0.8403 0.4000 0.8403
BTCRUB-BTCUSD 0.4000 0.8538 0.3333 0.8539 0.3500 0.8543
[…]
5.1.2 Grid Search and Reinforcement Learning
A grid search is used to find a set of profitable parameters, including the open/close threshold and window size during the training period (October 2023 to November 2023). Every iteration of the window-sliding pair trading experiments has one set of parameters (window size, open/close threshold) until exhaustion. We start off the exploration based on the experiential estimation of the asset characteristics. The most profitable parameter set will be used to test traditional pair trading during the test period (December 2023) and also for testing the proposed RL strategy.
The profitability in conducting a grid search is measured by the Total Compound Return (RTOT), where $V_{p}$ and $V_{p}^{\prime}$ are the values of the portfolio at the beginning of the period and the end of the period, and $t$ is the total length of the trading period in Equation (11):
$rtot=(V_{p}^{\prime}/V_{p})^{1/t}-1\times 100\%.$ (11)
During the training period, the most profitable parameter set is open threshold = 1.8 $z$-score, close threshold = 0.4 $z$-score, and window size = 900 intervals. Some example results of the grid search are presented in Table 3.
Table 3: Trading parameter tuning.
        OPEN _THRES         CLOS_THRES         PERIOD         RTOT (%)
        4.0         2.0         2000         0.0651
        4.0         0.5         500         0.5024
        3.0         1.0         500         0.9993
        3.0         0.5         1000         0.8932
        3.0         0.5         500         1.0704
        2.5         0.3         700         2.1542
        2.5         0.5         700         1.5667
        3.0         0.3         700         1.3160
        2.1         0.4         700         2.5633
        2.1         0.3         800         2.6916
        2.3         0.4         800         2.3096
        2.1         0.4         800         2.8202
        2.0         0.4         1000         2.7339
        2.0         0.4         900         3.0400
        1.9         0.3         900         2.8989
        1.9         0.4         900         3.1077
        1.8         0.4         900         3.0565
The setup of RL-based pair trading relies on these parameters. The window size decides the retrospective length of the spread, and the thresholds decide the zones. Algorithms such as PPO and A2C are applicable to both discrete and continuous action spaces. Some algorithms, e.g., DQN, can only be used in discrete space, and DDPG is only applicable in a continuous space. Therefore, we adopt PPO, DQN, and A2C in RL pair trading, which decide the timing, and PPO, A2C, and SAC in RL pair trading, which decide both the timing and investment quantity. The algorithms are adopted from the Baseline3 collection (Raffin et al., 2021).
5.1.3 Evaluation Metrics
Our main concern is the highest profitability in trading techniques. We care about the cumulative return, which is the total profit for the testing period, as well as the annualized return Compound Annual Growth Rate (CAGR). With $V(t_{0})$ as the initial state, $V(t_{n})$ as the final state, and $t_{n}-t_{0}$ as the period of trading in years, the CAGR is Equation (12):
$\operatorname{CAGR}\left(t_{0},t_{n}\right)=\left(\frac{V\left(t_{n}\right)}{V\left(t_{0}\right)}\right)^{\frac{1}{t_{n}-t_{0}}}-1.$ (12)
[…]
Sharpe $\text{Sharpe Ratio}=\frac{R_{p}-R_{f}}{\sigma_{p}}.$ (13)
We also care about the strategies’ activities, such as the order count and win/loss ratio. The indicators used for comparison are presented in Table 4.
Table 4: Descriptive table of evaluation metrics.
          Profitability Indicator           Description
          Cumulative Return           Profit achieved during trading period
          CAGR           Compound Annual Growth Rate
          Sharpe Ratio           Risk-adjusted returns ratio
          Activity Indicator           Description
          Total Action Count           Total orders executed
          Win/Loss Action Count           Number of winning/losing trades
          Win/Loss Action Ratio           Ratio of winning to losing trades
          Max Win/Loss Action           Maximum profit/loss per Action in Bitcoin
          Avg Action Profit/Loss           Average profit/loss per trade in Bitcoin
          Time in Market           Percentage of time invested in the market
          Risk Indicator           Description
          Volatility (ann.)           Annualized standard deviation of returns
          Skew           Asymmetry of returns distribution
          Kurtosis           “Tailedness” of returns distribution
5.2 Experimental Results
We present the profitability and risk results from our experiments along with the trading indicators.
5.2.1 Result Comparison
Our work is compared with standard pair trading (Section 2.1) and state-of-the-art pair trading techniques (Section 3.2).
Our results are presented in Table 5. The results display a positive return for the traditional pair trading technique Gatev et al. (2006). The algorithm A2C displays a positive return for RL pair trading techniques. However, the PPO, SAC, and DQN algorithms do not perform as well as A2C. If we view A2C as the chosen algorithm for pair trading, the results show a steady income from pair trading. The traditional pair trading of Gatev et al. (2006) displays a stable income compared to others due to its rule-based execution stability.
[…]
Table 5: Evaluation metrics comparison between trading techniques.
Gatev et al. (2006) Kim and Kim (2019) RL1 RL2
RL Algo. NA PPO A2C DQN PPO A2C DQN PPO A2C SAC
Profitability
Cumulative Return 8.33% --0.16% --35.16% --35.79% 1.89% 9.94% --31.99% --77.81% 31.53% --87.12%
CAGR 195.12% --2.19% --99.71% --99.75% 30.05% 278.72% --99.56% --100.00% 3974.65% --100.00%
Sharpe Ratio 25.91 --1.67 --2.04 --2.60 5.44 32.74 --8.77 --1.99 94.34 --1.93
Activities
Total Action Count 490 43 1248 1062 1304 249 879 3443 229 2798
Won Action Count 284 24 600 503 578 240 232 842 162 917
Lost Action Count 206 19 648 559 726 9 647 2601 67 1881
Win/Loss Action Ratio 1.38 1.26 0.93 0.90 0.80 26.67 0.36 0.32 2.42 0.49
Max Win Action (USD) 75.35 163.52 606.75 606.75 43.72 121.74 70.59 307.78 648.87 160.15
Max Loss Action (USD) --27.73 --187.86 --763.70 --553.25 --108.51 --21.33 --282.84 --389.22 --64.97 --1456.43
Avg Win Action Profit/Loss (USD) 14.50 41.72 38.68 37.47 5.51 17.77 11.30 15.88 90.94 8.03
Avg Loss Action Profit/Loss (USD) --2.90 --54.68 --58.75 --60.78 --3.28 --7.06 --24.95 --17.78 --21.00 --23.49
Risk
Volatility (ann.) 6.01% 3.93% 51.43% 40.43% 3.61% 6.30% 11.92% 53.04% 27.30% 54.66%
Skew 1840 --54 --358 --358 --874 2673 --1899 --374 4314 --3048
Kurtosis 48,145 135 4138 4201 133,603 114,944 54,808 12,987 254,283 138,851
{adjustwidth}
[…]
Behavior-wise, PPO, DQN, and SAC tend to conduct excessive transactions that are not profitable. On the contrary, A2C have fewer trades but higher profits on each trade. RL 2 pair trading shows further fewer total actions because of the adjusted position action, where we do not consider a position adjustment as one trade until it is closed. Apart from the result in Table 5, the portfolio growth trend with the best-performing RL algorithm agent is presented in Figure 6 (a comparison with Gatev et al. (2006) is provided in Figure 7a in the Appendix). Most of the pair trading experiments, including Gatev et al. (2006), RL1, and RL2, display a stable upturn, which is ideal from the perspective of pair trading. From the drawdown graphs, we can observe that RL 1 produces fewer drawdowns compared to the non-RL pair trading method from Gatev et al. (2006) and has a significantly higher win/loss action ratio due to differences in threshold settings. However, RL1’s cumulative profit is not consistently higher, and when transaction fees are zero, its cumulative profit is slightly lower than that of the Gatev et al. method. RL2 displays the strongest profitability, despite a lower win/loss action ratio, due to its progressive trading strategy. In general, all three pair trading methods show the ability to generate stable income in a volatile trading market.
a RL 1 Pair Trading (A2C).
b RL 2 Pair Trading (A2C).
Figure 6: Comparison of portfolio value trends for RL 1 and RL 2 Pair Trading (A2C).
5.2.2 Effect of Transaction Cost
The profitability of high-frequency trading techniques is always significantly impacted by transaction costs. Cryptocurrency exchanges normally provide a large-volume discount scheme. The Binance fee ranges from 0.02% to even 0%, depending on the volume and the holding of their membership token. Considering that the users of these techniques may benefit from different transaction fee tiers, we explore the trading techniques under different transaction cost tiers as well. With more exploration under 0.05%, 0.01%, and 0% transaction costs compared to the default 0.02% transaction cost, we can see the significant impact of decreasing the transaction cost. The participating approaches are adopted with the most profitable algorithm based on the backtesting result (Table 5). We can see that trading techniques generally perform better under lower transaction costs. RL-based techniques tend to perform more trades when the transaction costs are lower.
6 Discussion and Conclusions
Pair trading has been a popular algorithmic trading method for decades. The in-demand high-frequency trading domain requires a fast-track decision-making process. However, the traditional rule-based pair trading technique lacks the flexibility to cater to volatile market movement. In this research, we proposed a mechanism to adopt Reinforcement Learning (RL)to observe the market and produce profitable pair trading decisions. The first adoption of Reinforcement Learning pair trading grants the RL 1 agent the flexibility to decide action timing. The second adoption of Reinforcement Learning 2 pair trading further gives the RL agent the access to decide the timing and invest quantity.
We compared it to the traditional rule-based pair trading technique (Gatev et al., 2006) and a state-of-the-art RL pair trading technique (Kim and Kim, 2019) for December 2023 in the cryptocurrency market for BTCEUR and BTCGBP under a standard future 0.02% transaction cost. Kim and Kim’s method does not perform well in the cryptocurrency world. Gatev et al.’s method achieved 8.33% per trading period. Our first adoption of the RL 1 method achieved 9.94%, and the second adoption of the RL 2 method achieved 31.53% returns during the trading period. The outperformance is generally consistent across different transaction costs. The evaluation metrics show that RL-based techniques are generally more active than traditional techniques in the cryptocurrency market under various transaction costs. In general, our trading methods have greater market participation than Gatev et al.’s traditional rule-based pair trading and Kim and Kim’s threshold-adaptive RL pair trading (Tables 5 and 6).
Table 6: Evaluation metrics comparison under different transaction costs.
Indicators Trading Approaches
0.05% Transaction Fee Gatev et al. (2006) Kim and Kim (2019) RL1 RL2
Cumulative Profit 5.02% -0.26% 5.76% 7.40%
Sharpe Ratio 14.60 -2.34 21.00 7.82
Total Action Count 490 43 154 207
Won Action Count 246 23 152 110
Lost Action Count 244 20 2 97
Win/Loss Action Ratio 1.01 1.15 76.00 1.13
Max Win Action (USD) 72.82 114.76 43.36 606.22
Max Loss Action (USD) --30.26 --169.52 --8.30 --168.81
Avg Win Action Profit/Loss (USD) 13.57 37.94 16.10 70.99
Avg Loss Action Profit/Loss (USD) --4.99 --47.68 --5.64 --48.28
0.01% Transaction Fee Gatev et al. (2006) Kim and Kim (2019) RL1 RL2
Cumulative Profit 9.43% --1.13% 9.88% 33.99%
Sharpe Ratio 29.84 --7.07 33.24 104.40
Total Action Count 490 43 251 181
Won Action Count 317 20 242 149
Lost Action Count 173 23 9 32
Win/Loss Action Ratio 1.83 0.87 26.89 4.66
Max Win Action (USD) 76.20 65.99 121.74 675.23
Max Loss Action (USD) --26.88 --169.66 --21.33 --27.91
Avg Win Action Profit/Loss (USD) 13.93 24.88 17.52 98.74
Avg Loss Action Profit/Loss (USD) --2.48 --44.81 --7.06 --10.79
0% Transaction Fee Gatev et al. (2006) Kim and Kim (2019) RL1 RL2
Cumulative Profit 10.54% --2.00% 9.94% 80.92%
Sharpe Ratio 33.90 --5.76 32.74 2668.86
Total Action Count 483 43 249 429
Won Action Count 363 23 240 342
Lost Action Count 120 20 9 87
Win/Loss Action Ratio 3.02 1.15 26.67 3.93
Max Win Action (USD) 77.04 163.59 121.74 699.51
Max Loss Action (USD) --26.03 --217.54 --21.33 --72.21
Avg Win Action Profit/Loss (USD) 13.07 36.69 17.77 104.25
Avg Loss Action Profit/Loss (USD) --2.43 --80.76 --7.06 --16.68
The trading period is from 1 December 2023 to 31 December 2023 with an interest rate of 5.5%. RL 1 stands for the pair trading that allows Reinforcement Learning to decide upon the investment timing. RL 2 stands for dynamic scaling Reinforcement Learning pair trading that allows the RL agent to decide both investment timing and quantity. We adopted the PPO algorithm from Kim and Kim (2019) and A2C for RL1 and RL2.
Comparison between RL-based pair trading revealed the relationship between profitability and actions. Because financial trading is a special case of the RL environment, every action in financial trading is punished by the transaction cost. We notice that profitable RL trading often has a lower total trade count and higher profit per-win trade. That means the RL is better at spotting chances to make higher profits. RL 2 pair trading produces higher profits because of higher average wins from the position adjustment mechanism. When we adopt the righteous trading method, market volatility and transaction cost play crucial roles in profitable trading. Variable thresholds might not be adaptive to highly volatile markets, and fixed-threshold pair trading could lead to missing trading opportunities. RL with dynamic scaling investment could be a good direction in volatile market conditions if low transaction costs are achievable.
The techniques presented have certain limitations and offer opportunities for future work. One limitation is the relatively limited dataset scope, which could be expanded to include more diverse assets and longer timeframes to improve generalization. Additionally, focusing only on two-leg strategies restricts the potential for complex arbitrage opportunities; expanding to multi-leg strategies would enhance robustness. The computational demand during training can also be resource-intensive, requiring system parameter tuning. The model lacks consideration for transaction costs, which might impact real-world profitability. A lack of direct comparison with traditional models is another shortcoming. Future work could involve developing the Reinforcement Learning (RL) approach to multi-leg strategies, integrating pair formation into the trading process, cross-validating across different environments, and experimenting with alternative reward functions to improve decision-making and risk management.
[…]
Conceptualization, A.M. and H.Y.; methodology, A.M. and H.Y.; software, A.M. and H.Y.; validation, H.Y.; formal analysis, H.Y.; investigation, H.Y.; resources, A.M. and H.Y.; data curation, H.Y.; writing—original draft preparation, H.Y.; writing—review and editing, A.M.; visualization, H.Y.; supervision, A.M.; project administration, A.M.; funding acquisition, Not applicable. All authors have read and agreed to the published version of the manuscript.
[…]
This research received no external funding
[…]
Not applicable.
[…]
The original data presented in the study are openly available from Binance Exchange accessed on 8 November 2024 at (https://data.binance.vision/).
[…]
The authors declare no conflicts of interest.
[…]
Appendix A
a
b
Figure 7: Comparison of Pair Trading strategies from (a) Gatev et al. (2006) and (b) Kim and Kim (2019).
{adjustwidth}
[…]